If 5.50 gg of the unknown compound contained 0.183 molmol of CC and 0.367 molmol of HH, how many moles of oxygen, OO, were in the sample?

Express your answer to three significant figures and include the appropriate units.

.
Answer #1

Mole = Weight(gm)/Molecular weight

For Carbon:

0.183 = Weight/12

Weight of Carbon = 0.183 X 12 = 2.196  gm

Weight of Hydrogen = 0.367 X 1 = 0.367 gm

Total weight of hydrocarbon = 5.50 gm

weight of carbon and hydrogen = 2.196 + 0.367 = 2.563 gm

Weight of Oxygen = 5.50 - 2.563 = 2.937 gm

moles of Oxygen = 2.937/16 = 0.183